### v1

# A Trade-Off between Path Entanglement and Quantum Sensitivity

Abstract
Entanglement often increases quantum measurement schemes’ sensitivity. However, we find that in precision measurements with zero-mean Gaussian states, such as squeezed states, entanglement between different paths degrades measurement sensitivity. We prove an inverse relationship between entanglement entropy and sensitivity for measurements of single-mode phase shifts in multimode systems and for phase shifts on both modes in two-mode systems. In the two-mode case, which models devices such as interferometers, we find that entanglement strongly degrades differential phase sensitivity. Finally, we show that minimizing entanglement between paths maximizes the phase sensitivity of $N$-mode systems with zero-mean Gaussian state inputs.
Introduction. Entanglement is a radically nonclassical aspect of quantum mechanics [1, 2]. It is central to our modern understanding of theoretical phenomena such as the Unruh effect and black hole radiation [3, 4, 5, 6, 7, 8], and its presence is critical for a diverse set of important applications, including quantum computation [9], teleportation [10], and imaging [11]. Its ubiquitous uses throughout quantum mechanics have led to attempts to quantify entanglement as a resource [12, 13, 14]. One such quantifier is entanglement entropy, useful for pure bipartite systems as it can be straightforwardly calculated from the system’s density matrix [15, 16, 17].
Entanglement also plays a key role in quantum metrology, which studies how to use entangled states such as squeezed states or N00N states to make measurements with sensitivities impossible to achieve using comparable classical states [18, 19, 20, 21, 22]. It is of fundamental interest to determine precisely how entanglement contributes to measurement sensitivity.
We focus on entanglement in the context of measuring small phase shifts on zero-mean, multimode Gaussian states. Gaussian states are of interest experimentally, with many feasible schemes for their physical creation [23, 24, 25, 26], and also theoretically as a model for various continuous quantum phenomena [15, 17, 27]. In the setting of quantum metrology, interferometers often employ Gaussian states. Interferometers form the basis of precision measurements of gravitational waves, local gravity, and direct imaging of black holes, among others [28, 29, 30, 31]. In a typical interferometer, two input Gaussian states are first combined on a beam splitter and then propagate for a distance before recombining on a second beam splitter. The recombined states are then measured. Between the device’s beam splitters, the two paths correspond to distinct spatial modes, and each experiences a phase shift. The phase shifts are called a “common phase” if they have the same magnitude and sign and a “differential phase” if they have the same magnitude and opposite sign. When one interferometer input is a coherent state, entanglement in the form of squeezing increases its sensitivity to differential phase shifts [32].
We show that in the zero-mean case, the opposite is true for entanglement between modes, i.e. between paths. Specifically, a state’s quantum Fisher information (QFI) quantifies its ultimate sensitivity to a parametrized unitary, such as a phase shift [33, 34, 35, 36]. We show that the QFI of a two-mode Gaussian state varies inversely with the entanglement entropy between the two modes. Moreover, the differential phase QFI is hurt the most by path entanglement, while the common phase QFI is unique in escaping the trade-off, due to the symmetry of the common-phase unitary. An illustration of this trade-off is that the maximally path entangled Einstein-Podolsky-Rosen (EPR) state [15] has no sensitivity at all to differential phase shifts, which is evident in the lack of active path length control in superdense coding experiments [37].
We distinguish this work from that of refs. [38, 39], which find that entanglement does not improve QFI by analyzing certain specific non-Gaussian states. However, the unentangled and entangled states they consider have different numbers of modes, meaning they exist on different Hilbert spaces and transform under different unitaries. Moreover, their results are not quantified through entanglement entropy. This makes a true comparison untenable. Importantly, we arrive at a robust notion of trade-off between path entanglement and sensitivity by comparing Gaussian states on the same Hilbert space transforming under the same unitary.
Gaussian States and Quantum Fisher Information. An $N$-mode bosonic system is described by $N$ pairs of canonically conjugate operators [17, 40],
$\hat{\mathbf{q}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})^{T}:=(\hat{q}_{1},\dots,\hat{q}_{2N})^{T}.$ (1)
Setting =2Planck-constant-over-$\hbar=2$, the nonzero canonical commutation relations are $[\hat{x}_{a},\hat{p}_{a}]=2i$. We call the state zero-mean whenever $\langle\hat{x}_{a}\rangle=\langle\hat{p}_{a}\rangle=0$. In this paper, we will assume that indices $i,j,k$ range from 00 to $2N$ while indices $a,b$ range from 00 to $N$.
The covariance matrix $\mathbf{V}$ of a zero-mean state is defined as the symmetric matrix $V_{ij}:=\langle\{\hat{q}_{i},\hat{q}_{j}\}\rangle/2$. $V_{ii}$ is the variance of $\hat{q}_{i}$, and $V_{ij}$ is the covariance between $\hat{q}_{i}$ and $\hat{q}_{j}$. In this paper, we focus on zero-mean Gaussian states, which are completely characterized by their covariance matrix $\mathbf{V}$ [17, 40, 41]. This is equivalent to how multivariate Gaussian probability distributions are completely characterized by their means and covariances. In fact, as shown in Appendix B, a pure, zero-mean Gaussian state is uniquely specified among all quantum states by its covariance matrix and the zero-mean condition.
Our input state consists of a collection of independent 1-mode squeezed states with squeezing levels $r_{1},\dots,r_{N}$, characterized by the covariance matrix
$\mathbf{V}_{\text{in}}=\text{diag}\left(e^{-2r_{1}},e^{2r_{1}},\ldots e^{-2r_{N}},e^{2r_{N}}\right).$ (2)
As discussed in Appendix B, any pure, zero-mean Gaussian state can be generated by acting on a covariance matrix of the form of $\mathbf{V}_{\text{in}}$ by a passive transformation, i.e. a photon-number preserving transformation that takes Gaussian states to Gaussian states. Each passive transformation is represented by a symplectic orthogonal matrix $\mathbf{K}$. Acting on the input state represented by $\mathbf{V}_{\text{in}}$ with this transformation yields a new state with covariance matrix $\mathbf{V}$ satisfying
$\mathbf{V}=\mathbf{K}\mathbf{V}_{\text{in}}\mathbf{K}^{-1}.$ (3)
We note that any passive transformation can be expressed as a sequence of beam splitters and phase shifts [42, 43].
We are interested in the sensitivity of measuring small phase shifts on a zero-mean Gaussian state with covariance matrix $\mathbf{V}$. Specifically, assume each mode $a$ receives a phase shift of magnitude $g_{a}\phi$, and we want to measure $\phi$. For instance, we can take $g_{1}=-g_{2}=1$ to represent a differential phase shift between modes $1$ and $2$.
The state’s sensitivity to $\phi$ can be found through the quantum Fisher information (QFI), which describes how fast the state changes with respect to $\phi$ [36]. Assume that $\phi$ is tightly centered around 00, and let the QFI be $H$. The quantum Cramer-Rao bound states that the variance of a locally unbiased estimator $\phi_{\text{est}}$ of $\phi$ based on a single POVM measurement, the most general type of measurement allowed by quantum mechanics [16], must satisfy $\operatorname{Var}\phi_{\text{est}}\geq 1/\sqrt{H}$ [33, 34, 35, 36] (in the Bayesian framework, the LHS can also be turned into the mean-squared error in our estimate of $\phi$ through the van Trees inequality [44]). The upshot is that higher QFIs yield better sensitivity. As we derive in Appendix A using properties of symplectic matrices, the QFI of a pure, zero-mean Gaussian state rotating under the phase shifts $g_{a}\phi$ is
$\begin{split}&H=\frac{\text{Tr}\left((\mathbf{VG})^{2}-\mathbf{G}^{2}\right)}{2},\ \text{where}\\ &\mathbf{G}\coloneqq\text{diag}(g_{1},g_{1},g_{2},g_{2},\dots,g_{N},g_{N}).\end{split}$ (4)
For mixed states, the equality becomes an upper bound, meaning $H\leq\text{Tr}\left((\mathbf{VG})^{2}-\mathbf{G}^{2}\right)/2$. We will argue that entanglement between modes generically decreases the QFI.
Figure 1: The tradeoff between entanglement and QFI. (a) a collection of squeezed states are combined by an arbitrary multimode passive transformation $\mathbf{K}$, which in general entangles the modes. One mode then receives a phase shift $\phi$ and all modes are measured with some POVM. (b) A pair of modes are combined by an arbitrary multimode passive transformation $\mathbf{K}$, which generally entangles the modes. The first mode receives a phase shift $g_{1}\phi$, and the second receives a phase shift $g_{2}\phi$. (c) plots the scenario in (b) where the passive transformation corresponds to phase shifting one of the modes by $\theta$ and then mixing the two modes on a 50/50 beam splitter. $\theta$ controls the entanglement. As the entanglement increases, the common-phase QFI is unaffected, while the single and differential-phase QFIs decrease. This plot takes $r_{1}=r_{2}=1/2$.
Tradeoff in The Single Phase Shift Scenario. A challenge in formulating the trade-off between QFI and path entanglement for multimode systems is that multipartite entanglement is difficult to define [45]. The situation simplifies when there is a natural way to partition the system’s modes into two sets.
We begin by considering the case when only one mode out of $N$ receives a phase shift, as shown in fig. 1(a). Here, the natural bipartition consists of the phase-shifted mode and the rest of the system. For convenience, assume it is the first mode that receives the phase shift, so $g_{1}=1,g_{a}=0\ \forall a>1$. Defining $\mathbf{V}_{1}$ as the upper left $2\times 2$ block of $\mathbf{V}$, LABEL:QFI_general_exact becomes
$H=\frac{\|\mathbf{V}_{1}\|^{2}}{2}-1,$ (5)
where $\|\cdot\|^{2}$ denotes the Frobenius norm squared, the sum of the squares of the entries ([46], pg. 7). In particular,
$\|\mathbf{V}_{1}\|^{2}=\operatorname{Tr}(\mathbf{V}_{1})^{2}-2\,\text{det}\left(\mathbf{V}_{1}\right).$ (6)
The trace and the determinant can be given a physical meaning as follows. By definition, the mean photon number of the phase-shifted mode is $n_{1}=\langle\hat{x}^{2}_{1}+\hat{p}_{1}^{2}\rangle/4-1/2$, which is related to the trace of this mode’s covariance matrix as $\operatorname{Tr}(\mathbf{V}_{1})=4n_{1}+2$.
On the other hand, the determinant is related to the entanglement between the phase-shifted mode and the rest of the system. Let $\rho_{1}$ be the reduced state achieved from tracing out all but the phase-shifted mode, and $\mu=\operatorname{Tr}(\rho_{1}^{2})$ its purity. The purity of this one-mode state monotonically decreases with its Von Neumann entropy, $S$, the standard measure of entanglement of a bipartite system. This follows from $S=\frac{1-\mu}{2\mu}\ln\left(\frac{1+\mu}{1-\mu}\right)-\ln\left(\frac{2\mu}{1+\mu}\right)$ [41]. Furthermore, the purity is related to the determinant as $\text{det}(\mathbf{V}_{1})=1/\mu^{2}$ [41] [^1] The formula in the citation differs by factors of 1/2 because we use =2Planck-constant-over-$\hbar=2$ instead of their $\hbar=1$ convention..
Using the formulas for trace and determinant in eq. 6, our first main result follows from eq. 5:
$H=8n_{1}^{2}+8n_{1}-\left(\frac{1}{\mu^{2}}-1\right):=H_{\text{sqz}}\left(n_{1}\right)-\left(\frac{1}{\mu^{2}}-1\right),$ (7)
where $H_{\text{sqz}}(n):=8n^{2}+8n$ is the QFI of a single squeezed state of photon number $n$ receiving a phase shift of $\phi$ (See Theorem 1 or [35, 48]). Now, $11/\mu^{2}-$ 1 monotonically increases with the entanglement between the phase-shifted mode and the rest of the system, yielding the trade-off between entanglement and sensitivity.
Since $\mu\leq 1$, we also get $H\leq H_{\text{sqz}}\left(n_{1}\right)$. Equality holds only when the phase-shifted mode is completely unentangled with the rest of the system, meaning $\mu=1$. Then $H$ reduces to the QFI of a 1-mode squeezed state. This makes sense because, without entanglement, the rest of the system is entirely unaffected by the phase shift, so the situation reduces to a 1-mode setup. Then eq. 7 is saying that the larger the entanglement, the further we get from this optimum.
The upper bound continues to hold for mixed states, since LABEL:QFI_general_exact is still an upper bound for mixed states. For pure states, we also have the following lower bound.
$H=\frac{\|\mathbf{V}_{1}\|^{2}}{2}-1\geq\frac{\operatorname{Tr}(\mathbf{V}_{1})^{2}}{4}-1=4n_{1}^{2}+4n_{1}=\frac{H_{\text{sqz}}\left(n_{1}\right)}{2}.$ (8)
The inequality follows from the fact that $\det(\mathbf{V}_{1})\leq||\mathbf{V}_{1}||^{2}/2$, which can be shown using the Cauchy-Schwarz inequality. Note that this inequality is saturated exactly when $\mathbf{V}_{1}$ is proportional to the identity, or in other words, when it represents a thermal state. Hence, this lower bound is achieved when the phase-shifted mode is maximally entangled with the rest of the system, such as for an EPR state.
Putting our two bounds together, we see that for pure states, $H$ is constrained within a factor of $2$ as
$\frac{1}{2}H_{\text{sqz}}\left(n_{1}\right)\leq H\leq H_{\text{sqz}}\left(n_{1}\right).$ (9)
Tradeoff in the Two-Mode Scenario. Another system for which there is a natural bipartition is the two-mode system, as shown in fig. 1(b). In the previous section, we restricted ourselves to a phase shift on a single mode but allowed an arbitrary number of modes in the whole system. Here, we will restrict ourselves to two modes in the system but allow a phase shift on both modes.
With two modes, LABEL:QFI_general_exact yields (see Appendix A)
$H=\left(g_{1}^{2}-g_{1}g_{2}\right)H_{1}+\left(g_{2}^{2}-g_{1}g_{2}\right)H_{2}+g_{1}g_{2}H_{\text{com}},$ (10)
where we have defined $H_{a}$ for $a\in\{1,2\}$ to be the QFI from a phase shift on just the $a^{\text{th}}$ mode, i.e. $g_{a}=1$ and $g_{b\neq a}=0$. Similarly, $H_{\text{com}}$ is the QFI from a common phase shift with $g_{1}=g_{2}=1$. $H_{1}$ and $H_{2}$ obey eq. 7 and are therefore affected by the entanglement between the two modes.
On the other hand, using LABEL:QFI_general_exact, $H_{\text{com}}=\operatorname{Tr}(\mathbf{V}^{2})/2-2$. This is independent of entanglement: using eq. 3, we have $\operatorname{Tr}(\mathbf{V}_{\text{in}}^{2})=\operatorname{Tr}(\mathbf{V}^{2})$, so $H_{\text{com}}$ depends only on the initial squeezed states and not the entangling transformation $\mathbf{K}$. Moreover, we can compute $H_{\text{com}}$ of the initial input states by adding their QFIs, since the inputs are independent  [^2] This is a general fact (see eq.72 in [33]) that can be seen in our case from LABEL:QFI_general_exact: $H_{\text{com}}=H_{\text{sqz}}\left(n_{\text{in},1}\right)+H_{\text{sqz}}\left(n_{\text{in},2}\right)$. In summary, eq. 10 decomposes the QFI into the sum of three QFIs based on simpler scenarios.
We can make the trade-off between QFI and entanglement clearer by substituting eq. 7 into eq. 10 and using that the purity $\mu$ is the same regardless of which mode is traced out, by the Schmidt decomposition  [^3] A standard proof of the Schmidt decomposition from pg. 704-706 in [54] can be extended to an infinite-dimensional setting with a slight addition. The addition is to note that density operators are trace class and therefore compact ([55] Theorem VI.21), so a countable spectral decomposition can still be furnished for the reduced density operators ([55] Theorem VI.16). This yields
$\begin{split}H&=g_{1}g_{2}\left(H_{\text{sqz}}\left(n_{\text{in},1}\right)+H_{\text{sqz}}\left(n_{\text{in},2}\right)\right)\\ &\hskip 10.00002pt+\left(g_{1}^{2}-g_{1}g_{2}\right)H_{\text{sqz}}\left(n_{\text{rot},1}\right)\\ &\hskip 10.00002pt+\left(g_{2}^{2}-g_{1}g_{2}\right)H_{\text{sqz}}\left(n_{\text{rot},2}\right)-\left(g_{1}-g_{2}\right)^{2}\left(\frac{1}{\mu^{2}}-1\right),\end{split}$ (11)
where $n_{\text{in},a}$ is the initial mean photon number in mode $a$ and $n_{\text{rot},a}$ is the mean number of photons in mode $a$ right before the phase shifts. Recall that $11/\mu^{2}-$ 1 represents the entanglement between the two modes. Therefore, the sensitivity of a differential phase measurement, where $g_{1}=-g_{2}$, is most damaged by entanglement. The common phase measurement, $g_{1}=g_{2}$, is unique in escaping the trade-off between entanglement and sensitivity.
We also note that if the initial squeezed states are equally squeezed and hence have the same photon number, after the entangling transformation $\mathbf{K}$, their photon numbers will remain the same (see proposition 3 in section C.2). In other words, we can write $n_{\text{in},1}=n_{\text{rot},1}=n_{\text{in},2}=n_{\text{rot},2}:=n$, and eq. 11 simplifies to
$H=\left(g_{1}^{2}+g_{2}^{2}\right)H_{\text{sqz}}(n)-(g_{1}-g_{2})^{2}\left(\frac{1}{\mu^{2}}-1\right).$ (12)
Additionally, for the maximally path entangled EPR state, the lower bound from eq. 9 yields $21/\mu^{2}-1=H_{\text{sqz}}(n)/$ 2, so eq. 12 simplifies further to
$H=\frac{(g_{1}+g_{2})^{2}}{2}H_{\text{sqz}}(n).$ (13)
As a result, we see that EPR states are sensitive to common phase shifts but completely insensitive to differential phase shifts.
Figure 1(c) plots the sensitivity of such a pair of equally squeezed states to single, common, and differential phase shifts as their entanglement is varied. As expected, the common phase shift sensitivity is independent of entanglement. Meanwhile, when the state is maximally path entangled, the single mode phase shift drops to half its maximal value while the differential phase sensitivity vanishes.
Decoupled Squeezed States Optimize the QFI. In the $N$-mode case with general $g_{a}$ values, there are many relevant notions of path entanglement because there is no clear bipartition, making it difficult to write a formula for the QFI as a function of the path entanglement. However, there is always a well-defined notion of no entanglement: the state decomposes as a tensor product  [^4] In cases with identical particles, the definition of no entanglement can sometimes be more subtle due to the need to (anti-)symmetrize [56]. However, since we are employing a second quantization approach, these symmetrization issues do not arise. In particular, we are interested in entanglement between (spatial) modes, and these modes are concatenated straightforwardly through the tensor product.. We show that the QFI attains its maximum when path entanglement is absent.
To formalize this, we say $\mathbf{V}$ is decoupled when the covariances vanish between all pairs of modes. In our setup, this condition is equivalent to the state having no path entanglement:
Proposition 1.
$\mathbf{V}=\mathbf{K}\mathbf{V}_{\text{in}}\mathbf{K}^{-1}$ is decoupled if and only if the state is a tensor product of squeezed states with $\mathbf{V}_{\text{in}}$’s input squeezing levels, $r_{1}\dots r_{N}$, in some order and orientation.
For a proof, see Appendix C.1.
We say that a decoupled state is properly ordered if given any modes $a$ and $b$, if $g_{a}^{2}>g_{b}^{2}$, then mode $a$ is at least as squeezed as mode $b$. For convenience, let’s assume the input squeezing levels satisfy $r_{1}\geq\dots\geq r_{N}$ (since we can always permute the input states using $\mathbf{K}$), and label the modes such that the rotation rates satisfy $g_{1}^{2}\geq\ldots\geq g_{N}^{2}$. We have (see appendix C for the proof)
Theorem 1.
Optimal QFI states are exactly those that can be reached from a decoupled, properly ordered state by applying passive transformations within modes having the same $g_{a}$. These optimal QFI states have
$H=2\sum_{a=1}^{N}g_{a}^{2}\sinh^{2}\left(2r_{a}\right).$ (14)
For instance, since $\mathbf{V}_{\text{in}}$ is decoupled and properly ordered, if we do not act with a passive transformation (i.e. $\mathbf{K}=I$), we will have an optimal QFI state.
As a consequence of Theorem 1, the QFI of a pure, zero-mean Gaussian state passively transformed from 1-mode squeezed states with squeezing of at most $r$ is bounded by
$H\leq\|\mathbf{G}\|^{2}\sinh^{2}(2r).$ (15)
This bound demarcates the maximal phase sensitivity achievable with any passive symplectic transformation based on the initial collection of squeezed states.
A different QFI bound has been derived in ref. [52] by constraining the total mean photon number instead of the mean photon number per mode, as we do here. The total photon number constraint results in a bound with a similar form to ours, with the Schatten 2-norm of eq. 15 replaced with a Schatten-$\infty$ norm ([46], pg. 7). Experimental limitations tend to set a bound on the mean photon number per mode as opposed to the total mean photon number because the squeezing level per mode is constrained by loss and phase noise [53], whereas there is no similar experimental constraint on the total mean photon number in a multimode system.
We also note that eq. 15 generally will not hold if we remove the zero-mean condition. This is because the maximal squeezing level present in the initial squeezed states cannot be increased by a passive transformation, whereas if the initial states were displaced, passive transformations could increase their maximal displacement. As a result, with enough modes, it can become advantageous to inject displaced states instead of squeezed vacuum states, and the bound of eq. 15 can be evaded. Indeed, if we take $g_{1}=1$ and $g_{a}=0\ \forall a>1$, the RHS of eq. 15 is independent of $N$. However, if one allows a nonzero displacement in each mode, one can use $\mathbf{K}$ to siphon all the displacement into the phase-shifted mode, completely flouting the spatial mean photon number constraint and ensuring the QFI scales as $\sqrt{N}$. Then with enough modes, the displacement strategy will become more effective.
As another corollary of Theorem 1, since a 1-mode squeezed state is trivially decoupled and properly ordered, its QFI immediately follows as $H=2g_{1}^{2}\sinh^{2}(2r)$. Since for a 1-mode squeezed state, $n=\sinh^{2}(r)$, if we take $g_{1}=1$, we get the familiar $H_{\text{sqz}}(n)=8n^{2}+8n$. In general, letting $n_{\text{in},i}$ be the mean photon number in mode $i$, we could then write the optimal QFI as
$H=\sum_{a=1}^{N}g_{a}^{2}H_{\text{sqz}}\left(n_{\text{in},a}\right).$ (16)
This is just the sum of the QFIs from each mode of a decoupled, properly ordered state, which is expected since the QFI is additive across independent states (see eq.72 in [33]).
Finally, if all the $g_{a}$ are distinct, then the symplectic orthogonal transformations in Theorem 1 can only be phase shifts. Therefore, every optimal QFI state in this case must be decoupled.
Conclusion. In this paper, we elucidate a tradeoff between entanglement between the modes of a multimode Gaussian system and the sensitivity of the system to phase shifts. Specifically, we find that unless two modes receive the same phase shift, entanglement between these modes will degrade the QFI. We show this explicitly based on entanglement entropy for naturally bipartite systems, these being two-mode Gaussian states and $N$-mode Gaussian states with only one mode receiving a phase shift. Finally, in the general $N$-mode case, we describe the optimal QFI states and find that they indeed have no entanglement between modes receiving different phase shifts. The value of optimal QFI is of independent interest; keeping the $g_{a}$ fixed in eq. 16 and taking the limit where the squeezing ratios $r_{i}$ are large, this achieves Heisenberg scaling, since the QFI scales as the square of the photon number.
Acknowledgements.
We gratefully acknowledge the support of the National Science Foundation (NSF) through the LIGO operations cooperative agreement PHY18671764464 and NSF award 2308969.
Appendix A Details on the Setup and QFI Calculations
An $N$-mode bosonic system is described by $N$ pairs of canonically conjugate operators [17, 40]
$\hat{\mathbf{q}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})^{T}=(\hat{q}_{1},\dots,\hat{q}_{2N})^{T}.$ (17)
Setting =2Planck-constant-over-$\hbar=2$, the canonical commutation relations are $[\hat{q}_{j},\hat{q}_{k}]=2i\Omega_{jk}$, where $\mathbf{\Omega}$ is the matrix of the symplectic form:
$\mathbf{\Omega}:=\begin{bmatrix}\mathbf{\omega}&&\\ &\ddots&\\ &&\mathbf{\omega}\end{bmatrix}\quad\mathbf{\omega}:=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}.$ (18)
We assume the state is zero-mean, meaning $\langle\hat{q}_{i}\rangle=0$.
We are interested in the sensitivity of measuring small phase shifts on a zero-mean Gaussian state with covariance matrix $\mathbf{V}$. Define the generator
$\mathbf{g}=\begin{bmatrix}g_{1}\mathbf{\omega}&&\\ &\ddots&\\ &&g_{N}\mathbf{\omega}\end{bmatrix}.$ (19)
Here, $g_{a}\in\mathds{R}$ for $a=1,\dots,N$. Then,
$\mathbf{V}(\phi)=e^{\phi\mathbf{\mathbf{g}}}\mathbf{V}e^{-\phi\mathbf{\mathbf{g}}}$ (20)
is the covariance matrix of the state after a phase shift of magnitude $g_{a}\phi$ for each mode $a$.
The QFI of a pure, zero-mean Gaussian state is given by [57, 58, 52]
$H=\frac{1}{4}\text{Tr}\left(\mathbf{V}^{-1}\mathbf{V}^{\prime}\mathbf{V}^{-1}\mathbf{V}^{\prime}\right).$ (21)
where $\mathbf{V}^{\prime}=\partial_{\phi}\mathbf{V}(\phi)|_{\phi=0}$. By differentiating eq. 20, we find $\mathbf{V}^{\prime}=[\mathbf{g},\mathbf{V}]$, which allows us to expand eq. 21 into
$\frac{1}{2}\text{Tr}\left(\mathbf{g}^{2}-\mathbf{V}\mathbf{g}\mathbf{V}^{-1}\mathbf{g}\right).$ (22)
To handle $\mathbf{V}^{-1}$, note that since the state is pure, $\mathbf{V}$ is symmetric and symplectic by Proposition 2. Therefore, $\mathbf{V\Omega V}=\mathbf{\Omega}$, and since $\mathbf{\Omega}^{-1}=-\mathbf{\Omega}$, this implies
$\mathbf{V}^{-1}=-\mathbf{\Omega V\Omega}.$ (23)
Defining
$\begin{split}\mathbf{G}&\coloneqq\text{diag}(g_{1},g_{1},g_{2},g_{2},\dots,g_{N},g_{N})\\ &\hskip 2.16995pt=-\mathbf{\Omega}\mathbf{g},\end{split}$ (24)
we find that eq. 22 becomes
$H=\frac{\text{Tr}\left((\mathbf{VG})^{2}-\mathbf{G}^{2}\right)}{2}.$ (25)
We derived this for pure states, but for mixed states, the RHS is instead an upper bound. To see this, note that we can purify the mixed state into a pure, zero-mean Gaussian state by adding some extra modes [59], and we may assume these modes do not receive phase shifts. This can only increase the QFI since all measurements that could be performed before the addition can still be performed simply by ignoring the extra modes. However, the addition does not change the RHS of eq. 25 since $\mathbf{G}$ vanishes on the extra modes.
It is useful to express eq. 25 in terms of the blocks of $\mathbf{V}$. Given any $2N\times 2N$ matrix $\mathbf{M}$, we write $\mathbf{M}_{(a,b)}$ for the $2\times 2$ submatrix
$\mathbf{M}_{(a,b)}=\left[\begin{matrix}M_{2a-1,2b-1}&M_{2a-1,2b}\\ M_{2a,2b-1}&M_{2a,2b}\end{matrix}\right].$ (26)
Now, since $\mathbf{G}$ is diagonal, we have
$\begin{split}\operatorname{Tr}\left((\mathbf{V}\mathbf{G})^{2}\right)&=\sum_{a,b}\operatorname{Tr}(\mathbf{V}_{(a,b)}\mathbf{G}_{(b,b)}\mathbf{V}_{(b,a)}\mathbf{G}_{(a,a)})\\ &=\sum_{a,b}g_{a}g_{b}\operatorname{Tr}(\mathbf{V}_{(a,b)}\mathbf{V}_{(b,a)})\\ &=\sum_{a,b}g_{a}g_{b}\|\mathbf{V}_{(a,b)}\|^{2}.\end{split}$ (27)
In the last equality, we used that $\mathbf{V}_{(a,b)}=\mathbf{V}_{(b,a)}^{T}$ since $\mathbf{V}$ is symmetric. We can now express eq. 25 as
$H=\frac{1}{2}\sum_{a,b}g_{a}g_{b}\|\mathbf{V}_{(a,b)}\|^{2}-\sum_{a}g_{a}^{2}.$ (28)
For example, setting $g_{1}=1$, $g_{a}=0$ for $a>1$, we recover
$H=\frac{\|\mathbf{V}_{1}\|^{2}}{2}-1.$ (29)
In the two-mode case, we may rewrite eq. 28 as:
$\begin{split}H=\frac{g_{1}^{2}(\|\mathbf{V}_{(1,1)}\|^{2}-2)+g_{2}^{2}(\|\mathbf{V}_{(2,2)}\|^{2}-2)}{2}+g_{1}g_{2}\|\mathbf{V}_{(1,2)}\|^{2}\end{split}$ (30)
Next, setting $g_{1}=g_{2}=1$ in eq. 28, we have
$H_{\text{com}}=\frac{\|\mathbf{V}_{(1,1)}\|^{2}+\|\mathbf{V}_{(2,2)}\|^{2}+2\|\mathbf{V}_{(1,2)}\|^{2}}{2}-2.$ (31)
Using this to eliminate $\|\mathbf{V}_{(1,2)}\|$, eq. 30 becomes
$H=\left(g_{1}^{2}-g_{1}g_{2}\right)H_{1}+\left(g_{2}^{2}-g_{1}g_{2}\right)H_{2}+g_{1}g_{2}H_{\text{com}},$ (32)
where $H_{a}=\|\mathbf{V}_{(a,a)}\|^{2}/2-1$ is the QFI with $g_{a}=1$, $g_{b\neq a}=0$.
In the $N$-mode case with general $g_{a}$, we can similarly eliminate $\mathbf{V}_{(a,b)}$ with $a\neq b$ using the QFI $H_{\text{com};(a,b)}$ from $g_{a}=g_{b}=1$, $g_{c\neq a,b}=0$. This decomposes the QFI into the sum of terms corresponding to the QFI of a single phase shift and terms corresponding to the QFI of a common phase shift between any two of the modes:
$H=\sum_{a}g_{a}\left(g_{a}-\sum_{b\neq a}g_{b}\right)H_{a}+\sum_{a,b\ \text{with}\ a<b}g_{a}g_{b}H_{\text{com};a,b}$ (33)
However, each common phase shift term will no longer be independent of $\mathbf{K}$, instead also capturing a sense of entanglement between the two modes and the rest of the system. Furthermore, we cannot derive eq. 11 because the purities $\mu_{a}$ from tracing out all but the $a^{\text{th}}$ mode will no longer be the same. In other words, with $N$ modes and an arbitrary generator, there are too many relevant notions of entanglement to allow for simple formulas. This essentially boils down to the difficulty of precisely defining multipartite entanglement  [^5] One might also consider a scenario in which $g_{1},\dots g_{n}$ are 1 and $g_{m}=0$ for $m>n$; this would appear to be naturally bipartite between the first $i$ modes and the rest of the system. We leave it the future work as to whether a simple formula for the QFI in terms of an entanglement measure also exists in this case..
Appendix B Williamson and Euler Decompositions and Pure States
Given a positive definite matrix $\mathbf{V}$, the Williamson decomposition says that
$\mathbf{V}=\mathbf{S}\mathbf{V}_{\text{th}}\mathbf{S}^{T},$ (34)
where $\mathbf{V}_{\text{th}}=\text{diag}(\kappa_{1},\kappa_{1},\kappa_{2},\kappa_{2},\ldots,\kappa_{n},\kappa_{n})$ is the covariance matrix of a thermal state [17], and $\mathbf{S}$ is a symplectic matrix, meaning it preserves the symplectic form by satisfying $\mathbf{S}\mathbf{\Omega}\mathbf{S}^{T}=\mathbf{\Omega}$. The $\kappa_{a}$ are called the symplectic eigenvalues of $\mathbf{V}$. For an elegant proof, see [61].
Now, for convenience, we say that $\mathbf{V}$ is pure when its associated zero-mean Gaussian state is pure.
Proposition 2.
The following are equivalent:
  1. 1.
    $\mathbf{V}$ is pure.
  2. 2.
    $\det(\mathbf{V})=1$.
  3. 3.
    The symplectic eigenvalues of $\mathbf{V}$ are all 1.
  4. 4.
    $\mathbf{V}$ is symplectic.
  5. 5.
    The zero-mean Gaussian state with covariance matrix $\mathbf{V}$ is the unique zero-mean quantum state with covariance matrix $\mathbf{V}$.
Note that if one uses the $\hbar=1$ convention as in [41, 40, 62], then conditions 2-4 will need to be modified by suitable factors of 1/2.
Proof of Proposition 3.
We know that $\mathbf{V}$ is pure exactly when the associated Gaussian state has purity $\mu=1$. By [41], $\det(\mathbf{V})=1/\mu^{2}$, proving $(1)\iff(2)$. Every symplectic matrix satisfies $\det(\mathbf{S})=1$ (page 480 [62]), so $(4)\implies(2)$, and by eq. 34, $\det(\mathbf{V})=\det(\mathbf{V}_{\text{th}})=\prod_{a}\kappa_{a}^{2}$, so $(3)\implies(2)$.
Conversely, the generalized uncertainty principle yields $\kappa_{a}\geq 1$ [40, 62], so $(2)\implies(3)$. We note that often, the generalized uncertainty principle is stated as $\mathbf{V}+i\mathbf{\Omega}\geq 0$. This is equivalent since
$\mathbf{V}_{\text{th}}+i\mathbf{\Omega}=\mathbf{S}^{-1}(\mathbf{V}+i\mathbf{\Omega})(\mathbf{S}^{-1})^{T}\geq 0,$ (35)
where we used that $\mathbf{S}^{-1}$ is symplectic since the symplectic matrices form a group.
Next, $(3)$ implies $\mathbf{V}_{\text{th}}=\mathbf{I}$, reducing eq. 34 to $\mathbf{V}=\mathbf{S}\mathbf{S}^{T}$. The symplectic group is closed under the transpose, since inverting $\mathbf{S}^{-1}\mathbf{\Omega}(\mathbf{S}^{-1})^{T}=\mathbf{\Omega}$ yields $\mathbf{S}^{T}\mathbf{\Omega}^{-1}\mathbf{S}=\mathbf{\Omega}^{-1}$, and $\mathbf{\Omega}^{-1}=-\mathbf{\Omega}$. So $\mathbf{V}$ is the product of two symplectic matrices, meaning $(3)\implies(4)$.
So far, we have proven the equivalence of the first four statements. Notice that $(5)$ is true when $\mathbf{V}=\mathbf{I}$ is the identity, since a zero-mean state with identity covariance matrix has vanishing mean photon number and therefore must be the vacuum. If $(3)$ is true, then after the symplectic transformation $\mathbf{S}^{-1}$, the covariance matrix is $\mathbf{V}_{\text{th}}=\mathbf{I}$, so the resulting state is the vacuum. Reversing the transformations implies that the original state was Gaussian, proving $(3)\implies(5)$.
Conversely, we prove that if $(3)$ is false, $(5)$ is false (in fact, infinitely far off). If $\kappa_{a}>1$ for some $a$, it suffices to prove that the number of zero-mean states with covariance matrix $\mathbf{V}_{\text{th}}$ is infinite, since we can transform any of these into having the desired $\mathbf{V}$. Since $\mathbf{V}_{\text{th}}$ is diagonal, it then suffices to prove the infiniteness in the 1-mode case, where $\mathbf{V}_{\text{th}}=\kappa\mathbf{I}$ with $\kappa>1$. We can achieve this covariance matrix by perturbing the vacuum as $c_{1}\left|0\right\rangle+c_{2}\left|j\right\rangle$, where $\left|j\right\rangle$ is the number state with $j$ photons; for any $j$ large enough, it is straightforward to verify that there will be constant $c_{1}$ and $c_{2}$ yielding a zero-mean state with covariance matrix $\kappa\mathbf{I}$. ∎
Note that $(3)\implies(5)$ is a formal way of saying that minimum uncertainty states are Gaussian. Also, instead of referring to pure, zero-mean Gaussian states throughout the paper, we could have referred to zero-mean states with a pure $\mathbf{V}$, yielding another way of motivating the states we are primarily considering.
We can further decompose the Williamson decomposition: the Euler decomposition says that any symplectic matrix $\mathbf{S}$ can be written as
$\mathbf{S}=\mathbf{K}_{1}\mathbf{S}(\mathbf{r})\mathbf{K}_{2},$ (36)
where $\mathbf{K}_{1}$ and $\mathbf{K}_{2}$ are symplectic orthogonal and $\mathbf{S}(\mathbf{r})=\text{diag}(e^{-r_{1}},e^{r_{1}},\ldots,e^{-r_{N}},e^{r_{N}})$ [17, 62]. Combining this with the above proposition, $\mathbf{V}=\mathbf{S}\mathbf{S}^{T}=\mathbf{K}_{1}\mathbf{S}(\mathbf{r})^{2}\mathbf{K}_{1}^{-1}$, so taking $\mathbf{V}_{\text{in}}=\mathbf{S}(\mathbf{r})^{2}$ proves that any pure state’s covariance matrix, $\mathbf{V}$, can be decomposed as
$\mathbf{V}=\mathbf{K}\mathbf{V}_{\text{in}}\mathbf{K}^{-1}.$ (37)
Appendix C Proof of the Optimal QFI Theorem
We prove Theorem 1 and associated claims.
C.1 Proof of Proposition 1: Characterizing Unentangled Gaussian States
Recall that $\mathbf{V}$ is decoupled if all covariances between different modes vanish. We now prove
Proposition 1.
$\mathbf{V}=\mathbf{K}\mathbf{V}_{\text{in}}\mathbf{K}^{-1}$ is decoupled if and only if the state is a tensor product of squeezed states with $\mathbf{V}_{\text{in}}$’s input squeezing levels, $r_{1}\dots r_{N}$, in some order and orientation.
This proposition is important to understanding Theorem 1 because it implies that the decoupled condition indeed signals a lack of modal entanglement.
We prove the proposition through a more general result that will also be useful later. Specifically, for zero-mean Gaussian states, we claim there is no entanglement between modes with vanishing covariance. More precisely, we claim that if $\mathbf{V}$ is block-diagonal with arbitrary blocks $\mathbf{V}_{m}$, then the state decomposes as the tensor product of zero-mean Gaussian states with the $\mathbf{V}_{m}$ as covariance matrices. Moreover, if the state is pure, so are the $\mathbf{V}_{m}$.
It is useful to introduce the notation that for a pure $\mathbf{V}$, the associated zero-mean state is |$\ket{\mathbf{V}}$; Proposition 2 implies that |$\ket{\mathbf{V}}$ is well-defined and Gaussian. Our claim becomes
$\mathbf{V}=\bigoplus_{m}\mathbf{V}_{m}\implies\ket{\mathbf{V}}=\bigotimes_{m}\ket{\mathbf{V}_{m}}.$ (38)
Proof.
We have
$\det(\mathbf{V}_{m})\geq 1.$ (39)
To see this, if the purity of the reduced state is $\mu_{m}$, recall $\det(\mathbf{V}_{m})=1/\mu_{m}^{2}$, and $\mu_{m}\leq 1$. Next, using Proposition 2, we also have
$\prod_{m}\det(\mathbf{V}_{m})=\det(\mathbf{V})=1.$ (40)
Therefore, $\det(\mathbf{V}_{m})=1$, meaning |$\ket{\mathbf{V}_{m}}$ exists by Proposition 2. Then $\bigotimes_{m}\ket{\mathbf{V}_{m}}$ is a zero-mean state that has covariance matrix $\mathbf{V}$, so it must be |$\ket{\mathbf{V}}$. ∎
We can now quickly prove Proposition 1:
Proof.
If $\mathbf{V}$ is decoupled, $\mathbf{V}=\bigoplus_{a}\mathbf{V}_{a}$, with each $\mathbf{V}_{a}$ being $2\times 2$. Then the |$\ket{\mathbf{V}_{a}}$ in eq. 38 are all pure, 1-mode Gaussian states, which are by definition squeezed states. The claim about the squeezing levels follows from the fact that the eigenvalues of $\mathbf{V}$ are the union of the eigenvalues of the $\mathbf{V}_{a}$. The converse is easy since the tensor product clearly precludes correlations across modes. ∎
C.2 Proof of Photon Majorization Result
A key step in the proof of Theorem 1 involves using majorization to bound the mean photon numbers of decoupled, zero-mean states after a passive transformation. We begin with a brief overview of majorization. Majorization has found uses in subjects ranging from economics to quantum information science [12, 63, 64, 13].
Definition 1 (Majorization).
Given two sequences $\mathbf{x}=\{x_{a}\}_{a=1}^{N}$ and $\mathbf{y}=\{y_{a}\}_{a=1}^{N}$, let $\mathbf{x}^{\downarrow}=\{x^{\downarrow}_{i}\}_{a=1}^{N}$ denote $\mathbf{x}$ written in nondecreasing order, and similarly for $\mathbf{y}^{\downarrow}$. We say that $\mathbf{x}$ majorizes $\mathbf{y}$, written $\mathbf{x}\succ\mathbf{y}$, if for all $k<N$,
$\sum_{a=1}^{k}x_{a}^{\downarrow}\geq\sum_{a=1}^{k}y_{a}^{\downarrow}$ (41)
and
$\sum_{a=1}^{N}x_{a}=\sum_{a=1}^{N}y_{a}.$ (42)
Intuitively, if $\mathbf{x}\succ\mathbf{y}$, then $\mathbf{y}$ looks like an averaged version of $\mathbf{x}$. Sequences with more inequality across their values will majorize those with less inequality.
Proposition 3.
Start with an $N$-mode, zero-mean state (not necessarily pure or Gaussian) with a decoupled covariance matrix. Let $\mathbf{n}_{\text{in}}=\{n_{\text{in},a}\}_{a=1}^{N}$ be the mean photon numbers in each mode. Mean photon numbers of $\mathbf{n}_{\text{tar}}=\{n_{\text{tar},a}\}_{a=1}^{N}$ are possible after some passive transformation if and only if
$\mathbf{n}_{\text{in}}\succ\mathbf{n}_{\text{tar}}.$ (43)
The proposition formalizes the intuition that passive transformations average out mean photon numbers. The zero-mean condition is important; if we allow displacement in the initial states, passive transformations can siphon the displacement across all modes into one, violating the result.
We now prove the proposition. Let the original, decoupled covariance matrix be $\mathbf{V}_{\text{in}}$. Note that we are proving a general theorem, so $\mathbf{V}_{\text{in}}$ need not be as it was defined in eq. 2. However, it should be decoupled in that the covariances between different modes vanish. Suppose $\mathbf{K}$ characterizes the passive transformation we apply to $\mathbf{V}_{\text{in}}$, leading to the transformed covariance matrix $\mathbf{V}=\mathbf{K}\mathbf{V}_{\text{in}}\mathbf{K}^{-1}$.
Instead of the $2N\times 2N$ symplectic orthogonal matrix $\mathbf{K}$, it will be easier to work with its corresponding $N\times N$ unitary matrix $\mathbf{U}$ [40]. We state the correspondence using the $2\times 2$ block notation eq. 26: because $\mathbf{K}$ commutes with the symplectic form $\mathbf{\Omega}$ from eq. 18, each block of $\mathbf{K}$ must take the form
[…]
where each $U_{ab}$ is a complex number. It then follows from the orthogonality of $\mathbf{K}$ that the $U_{ab}$ must be the entries of a unitary matrix $\mathbf{U}$. (Physically, this correspondence means that in the Heisenberg picture, the passive transformation can be written as either $\hat{\mathbf{a}}\mapsto\mathbf{U}\hat{\mathbf{a}}$, where $\hat{\mathbf{a}}=(a_{1},\ldots,a_{N})^{T}$ is the vector of lowering operators, or $\hat{\mathbf{q}}\mapsto\mathbf{K}\hat{\mathbf{q}}$.)
Next, define a doubly stochastic matrix to be a square matrix with nonnegative entries such that each of its rows and columns sums to $1$. Let $\mathbf{D}$ be the matrix whose entries satisfy $D_{ab}=|U_{ab}|^{2}$. Since $\mathbf{U}$ is unitary, $\mathbf{U}\mathbf{U}^{\dagger}=\mathbf{U}^{\dagger}\mathbf{U}=\mathbf{I}$, from which one deduces that $\mathbf{D}$ is doubly stochastic.
The following fundamental theorem characterizes majorization in terms of doubly stochastic matrices (see [12] pg. 575, [63] pg.14, or [64] pg.33):
Theorem 2 (Hardy, Littlewood, Polya).
Two sequences $\mathbf{x},\mathbf{y}$ satisfy $\mathbf{x}\succ\mathbf{y}$ if and only if there exists a doubly stochastic matrix $\mathbf{D}$ such that $\mathbf{D}\mathbf{x}=\mathbf{y}$.
Using this, the following lemma suffices to prove Proposition 3:
Lemma 1.
$\mathbf{D}\mathbf{n}_{\text{in}}=\mathbf{n}_{\text{tar}}.$
Proof.
Letting $\mathbf{2}$ be the vector containing all 2s of length $N$, we see that $\mathbf{D2}=\mathbf{2}$. So the lemma is equivalent to $\mathbf{D}(4\mathbf{n}_{\text{in}}+\mathbf{2})=4\mathbf{n}_{\text{tar}}+\mathbf{2}.$ This is useful because
$\begin{split}&4\mathbf{n}_{\text{in}}+\mathbf{2}=\{\text{Tr}\left(\left(\mathbf{V}_{\text{in}}\right)_{\left(a,a\right)}\right)\}_{a=1}^{N},\\ &4\mathbf{n}_{\text{tar}}+\mathbf{2}=\{\text{Tr}\left(\mathbf{V}_{\left(a,a\right)}\right)\}_{a=1}^{N}.\end{split}$ (45)
Now we can prove the lemma directly:
$\begin{split}\text{Tr}\left(\mathbf{V}_{\left(a,a\right)}\right)&=\text{Tr}\left(\sum_{b=1}^{N}\mathbf{K}_{\left(a,b\right)}\left(\mathbf{V}_{\text{in}}\right)_{\left(b,b\right)}\left(\mathbf{K}^{-1}\right)_{\left(b,a\right)}\right)\\ &=\sum_{b=1}^{N}\text{Tr}\left(\mathbf{K}_{\left(a,b\right)}\left(\mathbf{V}_{\text{in}}\right)_{\left(b,b\right)}\left(\mathbf{K}_{\left(a,b\right)}\right)^{-1}\right)\\ &=\sum_{b=1}^{N}\left|U_{ab}\right|^{2}\text{Tr}\left(\left(\mathbf{V}_{\text{in}}\right)_{\left(b,b\right)}\right)\\ &=\sum_{b=1}^{N}D_{ab}\text{Tr}\left(\left(\mathbf{V}_{\text{in}}\right)_{\left(b,b\right)}\right).\end{split}$ (46)
C.3 Proof of Lemma 1
We need one more result that will help translate the majorization result of the previous section to a bound on the QFI.
Lemma 2.
Let $\{c_{a}\}_{a=1}^{N}$ be a non-increasing sequence. Suppose $\mathbf{n}=\{n_{a}\}_{a=1}^{N}$ and $\tilde{\mathbf{n}}=\{\tilde{n}_{a}\}_{a=1}^{N}$ are arbitrary sequences with $\mathbf{n}$ nonincreasing. If $\mathbf{n}\succ\tilde{\mathbf{n}}$, then
$\sum_{a=1}^{N}c_{a}n_{a}\geq\sum_{a=1}^{N}c_{a}\tilde{n}_{a}.$ (47)
Intuitively, averaging out the $\mathbf{n}$ can only prevent us from taking advantage of the larger $c_{a}$.
We now prove the lemma. Some technical issues arise because the $c_{a}$ may not be distinct; let us define $\{A_{m}\}_{m=1}^{M}$ with $A_{m}$ as the $m$th largest number among the set of $c_{a}$. In particular, the $A_{m}$ are all distinct. We also define $a(m)$ to be the largest $a$ for which $c_{a}=A_{m}$.
For any sequence $\mathbf{x}=\{x_{a}\}_{a=1}^{N}$ and integer $a=1,\dots,N$, we further define $s_{a}(\mathbf{x}):=\sum_{a=1}^{k}x_{a}^{\downarrow}$. With this notation, Definition 1 of $\mathbf{x}\succ\mathbf{y}$ can be written as $s_{a}(\mathbf{x})\geq s_{a}(\mathbf{y})$ for $a<N$ and $s_{N}(\mathbf{x})=s_{N}(\mathbf{y})$.
With this notation, we see that
$\sum_{a=1}^{N}c_{a}n_{a}=\sum_{m=1}^{M}\left(A_{m}-A_{m+1}\right)s_{a\left(m\right)}(\mathbf{n}),$ (48)
where we defined $A_{M+1}=0$. In particular,
$\sum_{a=1}^{N}c_{a}(n_{a}-\tilde{n}_{a})=\sum_{m=1}^{M}\left(A_{m}-A_{m+1}\right)(s_{a\left(m\right)}(\mathbf{n})-s_{a\left(m\right)}(\tilde{\mathbf{n}})).$ (49)
Since the RHS is positive, we are done.
We are now ready to prove Theorem 1, which we restate here for convenience.
Theorem 1.
Optimal QFI states are exactly those that can be reached from a decoupled, properly ordered state by applying passive transformations within modes having the same $g_{a}$. These optimal QFI states have
$H=2\sum_{a=1}^{N}g_{a}^{2}\sinh^{2}\left(2r_{a}\right).$ (50)
The proof proceeds in two steps. First, we show that the QFI $H$ satisfies
$H\leq 2\sum_{a=1}^{N}g_{a}^{2}\sinh^{2}\left(2r_{a}\right).$ (51)
Then we show the bound is saturated for the claimed states.
C.4 Proving the QFI Upper Bound
We bound the RHS of eq. 25 using that $\text{Tr}\left((\mathbf{VG})^{2}\right)\leq\text{Tr}\left(\mathbf{V}^{2}\mathbf{G}^{2}\right)$, valid because $\mathbf{V}$ and $\mathbf{G}$ are hermitian  [^6] There are several ways of proving this inequality. One way, suggested by [52], is to realize that it is equivalent to $\text{Tr}\left(\left(i\left(\mathbf{VG}-\mathbf{PV}\right)\right)^{2}\right)\geq 0$. And this latter inequality is true because $i\left(\mathbf{VG}-\mathbf{PV}\right)$ is hermitian Equality holds exactly when this operator is zero.:
$H\leq\frac{\text{Tr}\left(\mathbf{G}^{2}\left(\mathbf{V}^{2}-\mathbf{I}\right)\right)}{2}.$ (52)
Equality holds if and only if $[\mathbf{V},\mathbf{G}]=0$. Next, define
$\mathbf{\tilde{V}}:=\mathbf{V}^{2}=\mathbf{K}\mathbf{V}_{\text{in}}^{2}\mathbf{K}^{-1}.$ (53)
This is a valid covariance matrix for a zero-mean state |$\tilde{\ket{\mathbf{V}}}$ made by passively transforming 1-mode squeezed states with squeezing $2r_{1},2r_{2},...,2r_{N}$. Further defining
$\mathbf{\tilde{V}}_{a}=\left[\begin{matrix}\tilde{V}_{2a-1,2a-1}&\tilde{V}_{2a-1,2a}\\ \tilde{V}_{2a,2a-1}&\tilde{V}_{2a,2a}\end{matrix}\right],$ (54)
we can write eq. 52 as
$H\leq\frac{1}{2}\sum_{a=1}^{N}g_{a}^{2}\left(\text{Tr}\left(\mathbf{\tilde{V}}_{a}\right)-2\right).$ (55)
Letting $\tilde{n}_{a}$ be the mean photon number in mode $a$ of |$\tilde{\ket{\mathbf{V}}}$, recall that $\text{Tr}\left(\mathbf{\tilde{V}}_{a}\right)-2=4\tilde{n}_{a}$. Therefore,
$H\leq 2\sum_{a=1}^{N}g_{a}^{2}\tilde{n}_{a}.$ (56)
We will have proved eq. 51 if we can show
$2\sum_{a=1}^{N}g_{a}^{2}\tilde{n}_{a}\leq 2\sum_{a=1}^{N}g_{a}^{2}\sinh^{2}\left(2r_{a}\right).$ (57)
This is where majorization theory enters. Applying Proposition 3 to |$\tilde{\ket{\mathbf{V}}}$, the initial state is described by $\mathbf{V}_{\text{in}}^{2}$, which has squeezing levels $2r_{1},\dots,2r_{N}$, so the initial photon numbers are $\sinh^{2}(2\mathbf{r})=\{\sinh^{2}(2r_{i})\}_{a=1}^{N}$. Writing $\tilde{\mathbf{n}}=\{\tilde{n}_{a}\}_{a=1}^{N}$, the proposition yields
$\sinh^{2}(2\mathbf{r})\succ\tilde{\mathbf{n}}.$ (58)
Finally, recall that we ordered the $g_{a}^{2}$ and $r_{a}$ to be nonincreasing, so we can take $g_{a}^{2}=c_{a}$ and $\sinh^{2}(2\mathbf{r})=\mathbf{n}$ in Lemma 2, which proves eq. 57 and hence eq. 51.
In summary, we proved our general QFI bound through the following chain of (in)equalities:
$\begin{split}H&=\frac{\text{Tr}\left((\mathbf{VG})^{2}-\mathbf{G}^{2}\right)}{2}\leq 2\sum_{a=1}^{N}g_{a}^{2}\tilde{n}_{a}\leq 2\sum_{a=1}^{N}g_{a}^{2}\sinh^{2}\left(2r_{a}\right).\end{split}$ (59)
C.5 Optimal QFI States: Saturating the QFI Bound
Having proven the QFI bound eq. 51, we examine when the bound is saturated. To begin, the first inequality in eq. 59 is saturated exactly when $\mathbf{V}$ and $\mathbf{G}$ commute, i.e. when $\mathbf{V}$ is block-diagonal with one block for each unique $g_{a}$. If the blocks are $\mathbf{V}_{1},\mathbf{V}_{2},\ldots$, then we write $\mathbf{V}=\bigoplus_{m}\mathbf{V}_{m}$, where the index $m$ runs over the distinct blocks.
Now, the states claimed to be optimal in Theorem 1 have the right-block diagonal form to saturate the first inequality in eq. 59. Observe that properly ordered states also saturate the second inequality in eq. 59, so they are optimal QFI states. Furthermore, similar to the RHS of eq. 55, the quantity $2\sum_{a=1}^{N}g_{a}^{2}\tilde{n}_{a}$ can be rewritten in terms of $\text{Tr}(\mathbf{V}_{m}^{2})$, and these are invariant under symplectic orthogonal transformations (which just conjugate the blocks). Thus, the claimed states have the same QFI as a properly ordered state, so they are also optimal QFI states.
Next, we prove that these are the only states saturating the bound. Again, an optimal $\mathbf{V}$ must have the block-diagonal structure. Then by eq. 38 each of the $\mathbf{V}_{m}$ are covariance matrices of pure Gaussian states, so each must obey the decomposition eq. 37. Therefore, we can decouple each of the $\mathbf{V}_{m}$ using a symplectic orthogonal transformation, and again, $2\sum_{a=1}^{N}g_{a}^{2}\tilde{n}_{a}$ will be unchanged in the process. But once the entire $\mathbf{V}$ is decoupled, in order to saturate the second inequality in eq. 59, the state must be properly ordered; otherwise, we can increase $2\sum_{a=1}^{N}g_{a}^{2}\tilde{n}_{a}$ by switching the 1-mode squeezed states in modes $a$ and $b$ that violate proper ordering. This completes the proof of Theorem 1.